OpenStudy (anonymous):

Solve each of these systems using the substitution method. 2x + 3y = -4 x − 2y = 5 x = 1, y = -2 x = -1, y = 2 x = 1, y = 2 x = -2, y = -1

OpenStudy (darkigloo):

Isolate one of the variable in one equation. It's easier to do this with x-2y=5. \[x - 2y =5 \] \[x = 5 + 2y \] Now that you have x, plug it into the other eqation and solve for y.

OpenStudy (anonymous):

its 7

OpenStudy (darkigloo):

\[2x + 3y = -4 \] \[2(5 + 2y) + 3y = -4\] \[10 + 4y + 3y = -4\] \[10 + 7y = -4\] \[7y = -4 - 10\] \[7y = -14\] \[y = -2\] Do you get how I got -2?

OpenStudy (darkigloo):

To find x, just plug in x=-2 into x-2y=5.